OpenStudy (anonymous):

Jackson Middle School has 7 students in choir for every 1 student in band. If there are 55 students in band, how many students are in choir?

OpenStudy (anonymous):

all you do is divide 55 by 7

OpenStudy (anonymous):

what ever ur whole number is thats the answer

OpenStudy (anonymous):

1/7=55/x x=385

OpenStudy (anonymous):

so basically x = 55/7

OpenStudy (anonymous):

except basically thats completely wrong

OpenStudy (anonymous):

how